# Determinantal Point Process Attention Over Grid Cell Code Supports Out of Distribution Generalization

### 18 Jan 2024 v2

@@ Abstract
Deep neural networks have made tremendous gains in emulating human-like intelligence, and have been used increasingly as ways of understanding how the brain may solve the complex computational problems on which this relies. However, these still fall short of, and therefore fail to provide insight into how the brain supports strong forms of generalization of which humans are capable. One such case is out-of-distribution (OOD) generalization— successful performance on test examples that lie outside the distribution of the training set. Here, we identify properties of processing in the brain that may contribute to this ability. We describe a two-part algorithm that draws on specific features of neural computation to achieve OOD generalization, and provide a proof of concept by evaluating performance on two challenging cognitive tasks. First we draw on the fact that the mammalian brain represents metric spaces using grid cell code (e.g., in entorhinal cortex): abstract representations of relational structure, organized in recurring motifs that cover the representational space. Second, we propose an attentional mechanism that operates over the grid cell code using Determinantal Point Process (DPP), that we call DPP attention (DPP-A) - a transformation that ensures maximum sparseness in the coverage of that space. We show that a loss function that combines standard task-optimized error with DPP-A can exploit the recurring motifs in the grid cell code, and can be integrated with common architectures to achieve strong OOD generalization performance on analogy and arithmetic tasks. This provides both an interpretation of how the grid cell code in the mammalian brain may contribute to generalization performance, and at the same time a potential means for improving such capabilities in artificial neural networks.
@@ 1 Introduction
Abstract Representations of Relational Structure. The first component of the proposed framework relies on the idea that a key element underlying human-like OOD generalization is the use of low-dimensional representations that emphasize the relational structure between data points. Empirical evidence suggests that, for spatial information, this is accomplished in the brain by encoding the organism’s spatial position using a periodic code consisting of different frequencies and phases (akin to a Fourier transform of the space). Although grid cells were discovered for representations of space (Hafting et al., 2005; Sreenivasan & Fiete, 2011; Mathis et al., 2012) and used for guiding spatial behaviour (Erdem & Hasselmo, 2014; Bush et al., 2015), they have since been identified in non-spatial domains, such as auditory tones (Aronov et al., 2017), odor (Bao et al., 2019), episodic memory (Chandra et al., 2023), and conceptual dimensions (Constantinescu et al., 2016). These findings suggest that the coding scheme used by grid cells may serve as a general representation of metric structure that may be exploited for reasoning about the abstract conceptual dimensions required for higher level reasoning tasks, such as analogy and mathematics (McNamee et al., 2022). Of interest here, the periodic response function displayed by grid cells belonging to a particular frequency is invariant to translation by its period, and increasing the scale of a higher frequency response gives a lower frequency response and vice versa, making it invariant to scale across frequencies. This is particularly promising for prospects of OOD generalization: downstream systems that acquire parameters over a narrow training region may be able to successfully apply those parameters across transformations of translation or scale, given the shared structure (which can also be learned (Cueva & Wei, 2018; Banino et al., 2018; Whittington et al., 2020)).
@@
Figure 1: Schematic of the overall framework. Given a task (e.g., an analogy to solve), inputs (denoted as $\{A,B,C,D\}$) are represented by the grid cell code, consisting of units ("grid cells") representing different combinations of frequencies and phases. Grid cell embeddings ($\bm{x}_{A},\bm{x}_{B},\bm{x}_{C},\bm{x}_{D}$) are multiplied elementwise (represented as a hadamard product $\odot$) by a set of learned attention gates $\bm{g}$, then passed to a inference module $\bm{R}$. The attention gates $\bm{g}$ are optimized using $\mathcal{L}_{DPP}$, which encourages attention to grid cell embeddings that maximize the volume of the representational space. The inference module outputs a score for each candidate analogy (consisting of $A,B,C$ and a candidate answer choice $D$). The scores for all answer choices are passed through a softmax to generate an answer $\hat{y}$, which is compared against the target $y$ to generate the task loss $\mathcal{L}_{task}$.
@@ 2 Approach
Figure 1 illustrates the general framework. Task inputs, corresponding to points in a metric space, are represented as a set of grid cell embeddings $\bm{x}_{t=1..T}$, that are then passed to a inference module $\bm{R}$. The embedding of each input is represented by the pattern of activity of grid cells that respond selectively to different combinations of phases and frequencies. Attention over these is a learned gating $\bm{g}$ of the grid cells, the gated activations of which ($\bm{x}\odot\bm{g}$) are passed to the inference module ($\bm{R}$) . The parameterization of $\bm{g}$ and $\bm{R}$ are determined by backpropagation of the error signal obtained by two loss functions over the training set. Note that learning of parameter $\bm{g}$ occurs only over the training space and is not further modified during testing (i.e. over the test spaces). The first loss function, $\mathcal{L}_{DPP}$ favors attentional gatings over the grid cells that maximize the DPP-A objective; that is, the "volume" of the representational space covered by the attended grid cells. The second loss function, $\mathcal{L}_{task}$ is a standard task error term (e.g., the cross entropy of targets $y$ and task outputs $\hat{y}$ over the training set). We describe each of these components in the following sections.
@@ 2.2.1 Grid cell code
The spatial frequencies of grids ($F$) begin at a value of $0.0028*2\pi$. Wei et al. (2015) have shown that, to minimize the number of variables needed to represent an integer domain of size $S$, the firing rate widths and frequencies should scale geometrically in a range $(\sqrt{2}$, $\sqrt{e})$, closely matching empirically observed scaling in entorhinal cortex (Stensola et al., 2012). We choose a scaling factor of $\sqrt{2}$ to efficiently tile the space. One consequence of this efficiency is that the total number of discrete frequencies in entorhinal cortex is expected to be small. Empirically, it has been estimated to be between 8-12 (Moser et al., 2015)  [^4] It seems likely that the use of grid cell code for abstraction in human cognition requires a considerably greater number of states $S$ than that used by the rodent for sensory encoding. However, given exponential scaling, the total number of frequencies is expected to remain low, increasing as a logarithm of $S$.. Here, we choose $N_{f}=9$ (dimension of $F$) as the number of frequencies. $A$ refers to a particular location in a two dimensional space, and 100 offsets ($A_{offset}$) are used for each frequency to evenly cover a space of $1000\times 1000$ locations using 900 grid cells. These offsets represent different phases within each frequency and since there are 100 of them, $N_{p}=100$. Hence $N_{p}\times N_{f}=900$, which denotes the number of grid cells. Each point from the set of 2D points for the stimuli in a task (described in Section 2.1), was represented using the firing rate of 900 grid cells which constituted the grid cell embedding for that point to form the inputs to our model.
@@ 2.2.2 DPP-A
Formally, we treat obtaining $\mathcal{L}_{DPP}$ as an approximation of a determinantal point process (DPP). A DPP $\mathcal{P}$ defines a probability measure on all subsets of a set of items $\mathcal{\bm{U}}=\{1,2,...N\}$. For every $\bm{u}\subseteq\mathcal{\bm{U}}$, $P(\bm{u})\propto\det(\bm{V_{u}})$. Here $\bm{V}$ is a positive semidefinite covariance matrix and $\bm{V_{u}}=[V_{ij}]_{i,j\in\bm{u}}$ denotes the matrix $\bm{V}$ restricted to the entries indexed by elements of $\bm{u}$. The maximum a posteriori (MAP) problem $argmax_{\bm{u}}\det(\bm{V_{u}})$ is NP-hard (Ko et al., 1995). However $f(\bm{u})=\log(\det(\bm{V_{u}}))$ satisfies the property of a submodular function, and various algorithms exist for approximately maximizing them. One common way is to approximate this discrete optimization problem by replacing the discrete variables with continuous variables and extend the objective function to the continuous domain. Gillenwater et al. (2012) proposed a continuous extension that is efficiently computable and differentiable:
@@
to construct $\bm{V}$, where in our case $\bm{m}$ are the variances of the grid cell embeddings $\bm{S}$ computed over the training region $M$, $N$ is the number of grid cells and $w_{m},b$ are hyperparameters with values of 1 and 0.1 respectively. The dimensionality of $\bm{V}$ is $N_{f}N_{p}\times N_{f}N_{p}(900\times 900)$. $\bm{g_{f}}$ are the gates of the grid cells belonging to the $f$th frequency , so $\bm{g_{f}}=\bm{g}[fN_{p}:(f+1)N_{p}]$, where $N_{p}$ is the number of phases for each frequency. $\bm{V_{f}}=\bm{V}[fN_{p}:(f+1)N_{p},fN_{p}:(f+1)N_{p}]$ is the restriction of $\bm{V}$ to the grid cell embeddings for $f$th frequency, so it captured the covariance of the grid cell embeddings belonging to the $f$th frequency. $\sigma$ is sigmoid non-linearity applied to $\bm{g_{f}}$, defined as $\sigma(\bm{g_{f}})=1/(1+e^{-\bm{g_{f}}})$, so that their values are between 0 and 1. $diag(\sigma(\bm{g_{f}}))$ converts vector $\sigma(\bm{g_{f}})$ into a matrix with $\sigma(\bm{g_{f}})$ as the diagonal of the matrix and rest entries are zero, which is multiplied with $\bm{V-I}$, which results in elementwise multiplication of $\sigma(\bm{g_{f}})$ with the column vectors of $\bm{V-I}$ Equation 6 which involved summation of the logarithm of the determinant of the gated covariance matrix of grid cell embeddings within each frequency, over $N_{f}$ frequencies was used to compute the negative of $\mathcal{L}_{DPP}$. Maximizing $\hat{F}$ gave the approximate maximum within frequency log determinant for each frequency $f\in[1,N_{f}]$, which we denote for the $f$th frequency as $\hat{F}_{f}$. In the second step of the training procedure, we used the $f_{max_{DPP}}$ grid cell frequency, where $f_{max_{DPP}}=\operatorname*{arg\,max}_{f\in[1,N_{f}]}\hat{F}_{f}$. In other words, we used the grid cell embeddings for the frequency which had the maximum within-frequency log determinant at the end of the first step, which we find are best at capturing the relational structure across the training and testing data, thereby promoting out-of-distribution generalization. In this step, we trained the inference module minimizing $\mathcal{L}_{task}$ over $N_{E_{task}}$ epochs. More details can be found in Appendix 7.10.
@@ 2.2.3 Inference module
where $\bm{W_{ii}}$,$\bm{W_{hi}}$,$\bm{W_{if}}$,$\bm{W_{hf}}$,$\bm{W_{io}}$,$\bm{W_{ho}}$,$\bm{W_{ic}}$,$\bm{W_{hc}}$ are weight matrices and $\bm{b_{ii}},\bm{b_{hi}},\bm{b_{if}},\bm{b_{hf}},\bm{b_{io}},\bm{b_{ho}},\bm{b_{ic}},\bm{b_{hc}}$ are bias vectors. $\bm{h}[t]$ and $\bm{c}[t]$ are the hidden state and the cell state at time $t$ respectively. The hidden state for the last timestep was passed through a linear layer with a single output unit to generate a score for the candidate completions for each problem. We used a single layered LSTM of 512 hidden units, which corresponds to the size of the hidden state, cell state, bias vectors, and the weight matrices. The hidden and cell state of the LSTM was reinitialized at the start of each sequence for each candidate completion.
For $\bm{R}$ constructed using a transformer, we used the standard multi-head self attention ($MHSA$) mechanism followed by a multi-layered percepton ($MLP$) - a feedforward neural network with one hidden layer, with layer normalization (Ba et al., 2016) ($LN$) to transform $\bm{x_{R}}$ at each layer of the transformer, defined as below:
@@
The seven scores (one for the correct completion and for six foil completions) were normalized using a softmax function, such that higher score would correspond to higher probablity and vice versa and the probabilities sum to 1. The inference module was trained using the task specific cross entropy loss ($\mathcal{L}_{task}=\operatorname{cross-entropy}$) between the softmax-normalized scores and the index for the correct completion (target), defined as $-log(softmax($scores)[target]). It is worth noting that the properties of equivariance hold, since the probability distribution after applying softmax remains the same when the transformation (translation or scaling) is applied to the scores for each of the answer choices obtained from the output of the inference module, and when the same transformation is applied to the stimuli for the task and all the answer choices before presenting as input to the inference module to obtain the scores. We also tried formulating the tasks as regression problems, the details of which can be found in Appendix 7.7.
While our model is not construed to be as specific about the implementation of the $\bm{R}$ module, we assume that — as a standard deep learning component — it is likely to map onto neocortical structures that interact with the entorhinal cortex and, in particular, regions of the prefrontal-posterior parietal network widely believed to be involved in abstract relational processes (Waltz et al., 1999; Christoff et al., 2001; Knowlton et al., 2012; Summerfield et al., 2020). In particular, the role of prefrontal cortex in the encoding and active maintenance of abstract information needed for task performance (such as rules and relations) has often been modeled using gated recurrent networks, such as LSTMs (Frank et al., 2001; Braver & Cohen, 2000), and posterior parietal cortex has long be known to support “maps” that may provide an important substrate for computing complex relations (Summerfield et al., 2020).

### 23 Jan 2024 v3

@@ Abstract
Deep neural networks have made tremendous gains in emulating human-like intelligence, and have been used increasingly as ways of understanding how the brain may solve the complex computational problems on which this relies. However, these still fall short of, and therefore fail to provide insight into how the brain supports strong forms of generalization of which humans are capable. One such case is out-of-distribution (OOD) generalization— successful performance on test examples that lie outside the distribution of the training set. Here, we identify properties of processing in the brain that may contribute to this ability. We describe a two-part algorithm that draws on specific features of neural computation to achieve OOD generalization, and provide a proof of concept by evaluating performance on two challenging cognitive tasks. First we draw on the fact that the mammalian brain represents metric spaces using grid cell code (e.g., in the entorhinal cortex): abstract representations of relational structure, organized in recurring motifs that cover the representational space. Second, we propose an attentional mechanism that operates over the grid cell code using Determinantal Point Process (DPP), that we call DPP attention (DPP-A) - a transformation that ensures maximum sparseness in the coverage of that space. We show that a loss function that combines standard task-optimized error with DPP-A can exploit the recurring motifs in the grid cell code, and can be integrated with common architectures to achieve strong OOD generalization performance on analogy and arithmetic tasks. This provides both an interpretation of how the grid cell code in the mammalian brain may contribute to generalization performance, and at the same time a potential means for improving such capabilities in artificial neural networks.
@@ 1 Introduction
Abstract Representations of Relational Structure. The first component of the proposed framework relies on the idea that a key element underlying human-like OOD generalization is the use of low-dimensional representations that emphasize the relational structure between data points. Empirical evidence suggests that, for spatial information, this is accomplished in the brain by encoding the organism’s spatial position using a periodic code consisting of different frequencies and phases (akin to a Fourier transform of the space). Although grid cells were discovered for representations of space (Hafting et al., 2005; Sreenivasan & Fiete, 2011; Mathis et al., 2012) and used for guiding spatial behaviour (Erdem & Hasselmo, 2014; Bush et al., 2015), they have since been identified in non-spatial domains, such as auditory tones (Aronov et al., 2017), odor (Bao et al., 2019), episodic memory (Chandra et al., 2023), and conceptual dimensions (Constantinescu et al., 2016). These findings suggest that the coding scheme used by grid cells may serve as a general representation of metric structure that may be exploited for reasoning about the abstract conceptual dimensions required for higher-level reasoning tasks, such as analogy and mathematics (McNamee et al., 2022). Of interest here, the periodic response function displayed by grid cells belonging to a particular frequency is invariant to translation by its period, and increasing the scale of a higher frequency response gives a lower frequency response and vice versa, making it invariant to scale across frequencies. This is particularly promising for prospects of OOD generalization: downstream systems that acquire parameters over a narrow training region may be able to successfully apply those parameters across transformations of translation or scale, given the shared structure (which can also be learned (Cueva & Wei, 2018; Banino et al., 2018; Whittington et al., 2020)).
@@
Figure 1: Schematic of the overall framework. Given a task (e.g., an analogy to solve), inputs (denoted as $\{A,B,C,D\}$) are represented by the grid cell code, consisting of units ("grid cells") representing different combinations of frequencies and phases. Grid cell embeddings ($\bm{x}_{A},\bm{x}_{B},\bm{x}_{C},\bm{x}_{D}$) are multiplied elementwise (represented as a Hadamard product $\odot$) by a set of learned attention gates $\bm{g}$, then passed to the inference module $\bm{R}$. The attention gates $\bm{g}$ are optimized using $\mathcal{L}_{DPP}$, which encourages attention to grid cell embeddings that maximize the volume of the representational space. The inference module outputs a score for each candidate analogy (consisting of $A,B,C$ and a candidate answer choice $D$). The scores for all answer choices are passed through a softmax to generate an answer $\hat{y}$, which is compared against the target $y$ to generate the task loss $\mathcal{L}_{task}$.
@@ 2 Approach
Figure 1 illustrates the general framework. Task inputs, corresponding to points in a metric space, are represented as a set of grid cell embeddings $\bm{x}_{t=1..T}$, that are then passed to the inference module $\bm{R}$. The embedding of each input is represented by the pattern of activity of grid cells that respond selectively to different combinations of phases and frequencies. Attention over these is a learned gating $\bm{g}$ of the grid cells, the gated activations of which ($\bm{x}\odot\bm{g}$) are passed to the inference module ($\bm{R}$). The parameterization of $\bm{g}$ and $\bm{R}$ are determined by backpropagation of the error signal obtained by two loss functions over the training set. Note that learning of parameter $\bm{g}$ occurs only over the training space and is not further modified during testing (i.e. over the test spaces). The first loss function, $\mathcal{L}_{DPP}$ favors attentional gatings over the grid cells that maximize the DPP-A objective; that is, the "volume" of the representational space covered by the attended grid cells. The second loss function, $\mathcal{L}_{task}$ is a standard task error term (e.g., the cross entropy of targets $y$ and task outputs $\hat{y}$ over the training set). We describe each of these components in the following sections.
@@ 2.2.1 Grid cell code
The spatial frequencies of grids ($F$) begin at a value of $0.0028*2\pi$. Wei et al. (2015) have shown that, to minimize the number of variables needed to represent an integer domain of size $S$, the firing rate widths and frequencies should scale geometrically in a range $(\sqrt{2}$, $\sqrt{e})$, closely matching empirically observed scaling in entorhinal cortex (Stensola et al., 2012). We choose a scaling factor of $\sqrt{2}$ to efficiently tile the space. One consequence of this efficiency is that the total number of discrete frequencies in the entorhinal cortex is expected to be small. Empirically, it has been estimated to be between 8-12 (Moser et al., 2015)  [^4] It seems likely that the use of grid cell code for abstraction in human cognition requires a considerably greater number of states $S$ than that used by the rodent for sensory encoding. However, given exponential scaling, the total number of frequencies is expected to remain low, increasing as a logarithm of $S$.. Here, we choose $N_{f}=9$ (dimension of $F$) as the number of frequencies. $A$ refers to a particular location in a two dimensional space, and 100 offsets ($A_{offset}$) are used for each frequency to evenly cover a space of $1000\times 1000$ locations using 900 grid cells. These offsets represent different phases within each frequency and since there are 100 of them, $N_{p}=100$. Hence $N_{p}\times N_{f}=900$, which denotes the number of grid cells. Each point from the set of 2D points for the stimuli in a task (described in Section 2.1), was represented using the firing rate of 900 grid cells which constituted the grid cell embedding for that point to form the inputs to our model.
@@ 2.2.2 DPP-A
Formally, we treat obtaining $\mathcal{L}_{DPP}$ as an approximation of a determinantal point process (DPP). A DPP $\mathcal{P}$ defines a probability measure on all subsets of a set of items $\mathcal{\bm{U}}=\{1,2,...N\}$. For every $\bm{u}\subseteq\mathcal{\bm{U}}$, $P(\bm{u})\propto\det(\bm{V_{u}})$. Here $\bm{V}$ is a positive semidefinite covariance matrix and $\bm{V_{u}}=[V_{ij}]_{i,j\in\bm{u}}$ denotes the matrix $\bm{V}$ restricted to the entries indexed by elements of $\bm{u}$. The maximum a posteriori (MAP) problem $argmax_{\bm{u}}\det(\bm{V_{u}})$ is NP-hard (Ko et al., 1995). However $f(\bm{u})=\log(\det(\bm{V_{u}}))$ satisfies the property of a submodular function, and various algorithms exist for approximately maximizing them. One common way is to approximate this discrete optimization problem by replacing the discrete variables with continuous variables and extending the objective function to the continuous domain. Gillenwater et al. (2012) proposed a continuous extension that is efficiently computable and differentiable:
@@
to construct $\bm{V}$, where in our case $\bm{m}$ are the variances of the grid cell embeddings $\bm{S}$ computed over the training region $M$, $N$ is the number of grid cells and $w_{m},b$ are hyperparameters with values of 1 and 0.1 respectively. The dimensionality of $\bm{V}$ is $N_{f}N_{p}\times N_{f}N_{p}(900\times 900)$. $\bm{g_{f}}$ are the gates of the grid cells belonging to the $f$th frequency , so $\bm{g_{f}}=\bm{g}[fN_{p}:(f+1)N_{p}]$, where $N_{p}$ is the number of phases for each frequency. $\bm{V_{f}}=\bm{V}[fN_{p}:(f+1)N_{p},fN_{p}:(f+1)N_{p}]$ is the restriction of $\bm{V}$ to the grid cell embeddings for $f$th frequency, so it captured the covariance of the grid cell embeddings belonging to the $f$th frequency. $\sigma$ is sigmoid non-linearity applied to $\bm{g_{f}}$, defined as $\sigma(\bm{g_{f}})=1/(1+e^{-\bm{g_{f}}})$, so that their values are between 0 and 1. $diag(\sigma(\bm{g_{f}}))$ converts vector $\sigma(\bm{g_{f}})$ into a matrix with $\sigma(\bm{g_{f}})$ as the diagonal of the matrix and the rest entries are zero, which is multiplied with $\bm{V-I}$, which results in elementwise multiplication of $\sigma(\bm{g_{f}})$ with the column vectors of $\bm{V-I}$. Equation 6 which involved summation of the logarithm of the determinant of the gated covariance matrix of grid cell embeddings within each frequency, over $N_{f}$ frequencies was used to compute the negative of $\mathcal{L}_{DPP}$. Maximizing $\hat{F}$ gave the approximate maximum within frequency log determinant for each frequency $f\in[1,N_{f}]$, which we denote for the $f$th frequency as $\hat{F}_{f}$. In the second step of the training procedure, we used the $f_{max_{DPP}}$ grid cell frequency, where $f_{max_{DPP}}=\operatorname*{arg\,max}_{f\in[1,N_{f}]}\hat{F}_{f}$. In other words, we used the grid cell embeddings for the frequency which had the maximum within-frequency log determinant at the end of the first step, which we find are best at capturing the relational structure across the training and testing data, thereby promoting out-of-distribution generalization. In this step, we trained the inference module minimizing $\mathcal{L}_{task}$ over $N_{E_{task}}$ epochs. More details can be found in Appendix 7.10.
@@ 2.2.3 Inference module
where $\bm{W_{ii}}$,$\bm{W_{hi}}$,$\bm{W_{if}}$,$\bm{W_{hf}}$,$\bm{W_{io}}$,$\bm{W_{ho}}$,$\bm{W_{ic}}$,$\bm{W_{hc}}$ are weight matrices and $\bm{b_{ii}},\bm{b_{hi}},\bm{b_{if}},\bm{b_{hf}},\bm{b_{io}},\bm{b_{ho}},\bm{b_{ic}},\bm{b_{hc}}$ are bias vectors. $\bm{h}[t]$ and $\bm{c}[t]$ are the hidden state and the cell state at time $t$ respectively. The hidden state for the last timestep was passed through a linear layer with a single output unit to generate a score for the candidate completions for each problem. We used a single layered LSTM of 512 hidden units, which corresponds to the size of the hidden state, cell state, bias vectors, and weight matrices. The hidden and cell state of the LSTM was reinitialized at the start of each sequence for each candidate completion.
For $\bm{R}$ constructed using a transformer, we used the standard multi-head self attention ($MHSA$) mechanism followed by a multi-layered perceptron ($MLP$) - a feedforward neural network with one hidden layer, with layer normalization (Ba et al., 2016) ($LN$) to transform $\bm{x_{R}}$ at each layer of the transformer, defined as below:
@@
The seven scores (one for the correct completion and for six foil completions) were normalized using a softmax function, such that a higher score would correspond to a higher probability and vice versa, and the probabilities sum to 1. The inference module was trained using the task specific cross entropy loss ($\mathcal{L}_{task}=\operatorname{cross-entropy}$) between the softmax-normalized scores and the index for the correct completion (target), defined as $-log(softmax($scores)[target]). It is worth noting that the properties of equivariance hold, since the probability distribution after applying softmax remains the same when the transformation (translation or scaling) is applied to the scores for each of the answer choices obtained from the output of the inference module, and when the same transformation is applied to the stimuli for the task and all the answer choices before presenting as input to the inference module to obtain the scores. We also tried formulating the tasks as regression problems, the details of which can be found in Appendix 7.7.
While our model is not construed to be as specific about the implementation of the $\bm{R}$ module, we assume that — as a standard deep learning component — it is likely to map onto neocortical structures that interact with the entorhinal cortex and, in particular, regions of the prefrontal-posterior parietal network widely believed to be involved in abstract relational processes (Waltz et al., 1999; Christoff et al., 2001; Knowlton et al., 2012; Summerfield et al., 2020). In particular, the role of the prefrontal cortex in the encoding and active maintenance of abstract information needed for task performance (such as rules and relations) has often been modeled using gated recurrent networks, such as LSTMs (Frank et al., 2001; Braver & Cohen, 2000), and the posterior parietal cortex has long been known to support “maps” that may provide an important substrate for computing complex relations (Summerfield et al., 2020).